### v1

@@ 10.12  Averages
$P(n)=\prod_{k=1}^{K}\frac{\left[(\Gamma+b_{k})T_{k}\right]^{n_{k}}}{n_{k}!}e^{-(\Gamma+b_{k})T_{k}},$ (90)
@@ 11 Conclusions
The original Garwood interval ticks the boxes on almost all of our desirable properties, including continuous monotonic and strictly nested behaviors and sensible $p$ values. The issue with it is that the interval lengths and over-coverage are larger than achievable if we relax these features. This fault is what led to the significant literature looking for “improvements”, generally without considering some of the possible drawbacks. Unfortunately, we know of no improvement that has the other desirable properties.
@@
It may be debated whether the properties such as nesting, continuity, and provision of sensible $p$ values are requirements for a confidence interval. At least superficially, several of our intervals provide shorter length and smaller overcoverage than the Garwood interval. If they are only used as descriptive intervals, do we care about bizzare behavior in other respects? We contend that the interval must at least include the maximum likelihood estimator, and we have seen that this rules out Crow&Gardner, and possibly others. But there are intervals besides Garwood, such as the Sterne and the likelihood ratio that do contain the MLE. They also seem to have reasonable scale properties (excluding FC). As long as we never look too closely, maybe this is ok?
@@
Unfortunately, as soon as we try to look at the confidence intervals for our observation at different confidence levels, as we sometimes do, we get into trouble. Intervals that change drastically for a small change in confidence level (e.g., Fig. 8b) lose their descriptive appeal. Of the intervals studied, the only two-sided exact interval that displays continuous dependence on confidence level is the Garwood. A similar issue at the other extreme is that of nesting. When we look at the intervals corresponding to two different confidence levels, we naturally expect the interval with the smaller confidence level to be a proper subset of the interval with a higher confidence level. Again, our only success among two-sided exact intervals is the Garwood interval.
@@
As noted in Sec. 10.11, properties of nesting and continuity are related to bimonoticity in the $p$ values (as a funtcion of the null hypothesis). Requiring the confidence intervals and the $p$ values to be part of a consistent framework is desirable for a consistent description of the result of the measurement. For example, a confidence interval that is closer to the null hypothesis should correspond to a higher $p$ value than an interval which is farther. We thus consider it important that our intervals provide sensible $p$ values. In particular, the $p$ values should exhibit a continuous bimonotonic dependence on the null hypothesis. Among two-sided exact intervals only the Garwood interval achieves this, and in fact, gives the two-sided $p$ values that we naively calculate (“propability of obtaining at result as different or more from the null hypothesis as observed”).
@@
While coverage and length are important, so are these other properties, and ensuring them seems worth erring on the conservative side in coverage. It is thus our recommendation to use the Garwood interval for the Poisson distribution. Fortunately, it is among the easiest to compute given an observation. As we have noted, it is also the default already for MATLAB, R, and perhaps other toolkits.
@@
We have cautioned that relying only on stated confidence intervals in averaging result of different measurements can be dangerous, and one should pay attention to the original sampling distribution when it is not normal.
@@
One of the important uses of confidence intervals is in the graphical representation of a measurement, especially when histogramming data where each bin is a sampling from an independent Poisson distribution. If backgrounds are subtracted, don’t be afraid to plot error bars that go negative.
We have examined a specific situation, sampling from a Poisson distribution with known background rate. Among several methods for computing exact confidence intervals as descriptive statistics, we recommend using the Garwood interval. We intend our discussion to be informative more generally. For example, if there are unknown nuisance parameters or the sampling distribution is not Poisson, the considerations here continue to apply in choosing methodology for descriptive statistics.

### v2

@@ 11.12 Averages
$\hbox{Prob}(n)=\prod_{k=1}^{K}\frac{\left[(\Gamma+b_{k})T_{k}\right]^{n_{k}}}{n_{k}!}e^{-(\Gamma+b_{k})T_{k}},$ (88)
@@ 12 Conclusions
We are ready to take our observations and come to some recommendations for descriptive intervals. We are insisting on exact confidence intervals, satisfying Eq. 6, as discussed in section 4.1. We are also requiring that our intervals be connected, as discussed in section 4.2. Describing the observation with $\sqrt{n}$ is commonly done and we have an intuitive feeling for what it means. However, it does not provide a probabilistic interpretation that can be used in functional dependences as discussed in sections 1 and 7. While $\sqrt{n}$ is useful and expedient, we look here for exact intervals. The Bayes’ intervals studied in section 8 do not satisfy Eq. 6 and in any event are not motivated as descriptive.
We have already argued that one-sided intervals (limits) are not as informative as two-sided intervals (section 5). The motivation for quoting a one-sided interval is typically to make a statement about the true value of the parameter. In this case, a Bayesian degree-of-belief analysis is recommended. If one insists on quoting a frequentist one-sided interval, then the “standard” methodology of section 5 is recommended. The $CL_{s}$ method (section 9.1) also produces an exact upper limit satisfying Eq. 6. However, it substantially overcovers for small $\theta$, with large intervals for small observations (Fig. 15), hence is less optimal compared with the standard limit. This is understandable because the $CL_{s}$ method is motivated with consideration about truth. It thus does not provide a good descriptive statistic.
The FC interval and the cut LR interval (or variations in section 9.2), while two-sided and exact, have restrictions on the interval in the case when $b\neq 0$. This restriction leads to intervals that may be very small, much smaller than the expected intuitive scale, when downward fluctuations occur. Hence the interpretation as description is clouded. While we note that FC [13] suggest to also quote the “sensitivity” of the experiment when this occurs, we conclude that there are better choices for descriptive intervals that do not have this feature.
We turn now to the conventional methods (section 6), giving two-sided exact intervals, which are of most interest as descriptive intervals. No choice stands out on all of the hoped-for features. The comparison here is more nuanced and some subjective choices must be made on which properties to emphasize over others.
@@
The original Garwood interval ticks the boxes on almost all of our desirable properties, including continuous monotonic and strictly nested behaviors and sensible $p$-values. The issue with it is that the interval lengths and over-coverage are larger than achievable if we relax these features. This fault is what led to the significant literature looking for “improvements”, generally without considering some of the possible drawbacks. Unfortunately, we know of no improvement that has the other desirable properties. Among intervals that do possess strict nestedness, the Garwood intervals are optimal as shortest length [1].
@@
It may be debated whether the properties such as nesting, continuity, and provision of sensible $p$-values are requirements for a confidence interval. Several of our intervals provide shorter length and smaller overcoverage than the Garwood interval. If they are only used as descriptive intervals, do we care about bizarre behavior in other respects? We contend that the interval must at least include the maximum likelihood estimator (and we insist here that for descriptive purposes the likelihood function should be evaluated without regard to knowledge of “truth”), and we have seen that this argues against Crow&Gardner (if low CL is considered), and possibly others. But there are intervals besides Garwood, such as the Sterne and the likelihood ratio that do contain the MLE. They also seem to have reasonable scale properties. As long as we never look too closely, maybe this is ok?
@@
Unfortunately, as soon as we try to look at the confidence intervals for our observation at different confidence levels, as we sometimes do, we get into trouble. Intervals that change drastically for a small change in confidence level (e.g., Fig. 8) lose their descriptive appeal. Of the intervals studied, the only two-sided exact interval that displays continuous dependence on confidence level is the Garwood. Another concern is that of nesting. When we look at the intervals corresponding to two different confidence levels, we naturally expect the interval with the smaller confidence level to be a proper subset of the interval with a higher confidence level. Again, our only success among two-sided exact intervals is the Garwood interval.
@@
As noted in section 11.11, properties of nesting and continuity are related to bimonotonicity in the $p$-values (as a function of the null hypothesis). Requiring the confidence intervals and the $p$-values to be part of a consistent framework is desirable for a consistent description of the result of the measurement. For example, a confidence interval that is closer to the null hypothesis should correspond to a higher $p$-value than an interval which is farther. We thus consider it important that our intervals provide sensible $p$-values. In particular, the $p$-values should exhibit a continuous bimonotonic dependence on the null hypothesis. Among two-sided exact intervals only the Garwood interval achieves this, and in fact, gives the two-sided $p$-values that we naively calculate (“probability of obtaining at result as different or more from the null hypothesis as observed”).
@@
While coverage and length are important, so are these other properties, and ensuring them seems worth erring on the conservative side in coverage. It is thus our recommendation to use the Garwood interval for the Poisson distribution. It appears to be the best that we can do while satisfying the intuitively hoped-for behavior. Fortunately, it is among the easiest to compute given an observation. As we have noted, it is also the default already for MATLAB, R, and perhaps other toolkits.
@@
We have cautioned that relying only on stated confidence intervals in averaging result of different measurements may be unjustified, and one should pay attention to the original sampling distribution when it is not normal.
@@
One of the important uses of confidence intervals is in the graphical representation of a measurement, especially when histogramming data where each bin is a sampling from an independent Poisson distribution. If backgrounds are subtracted, do not be afraid to plot error bars that go negative, whether using Garwood or $\sqrt{n}$. Given that there are multiple choices in use, it is helpful to say how the error bars on histograms are computed.
We have examined a specific situation, sampling from a Poisson distribution with known background rate. Among several methods for computing exact confidence intervals as descriptive statistics, we recommend using the Garwood interval. Our discussion has been deliberately restricted to a “simple case”, already with complexity and controversy. A similar discussion, with less complexity, for the normal distribution is in [24]. It may be remarked that the methodologies of Garwood, Sterne, Crow&Gardner, likelihood ratio, and score presented here, if applied to the normal case, all give the identical, familiar, results for normal two-sided confidence intervals.
More generally, there may be unknown nuisance parameters that are estimated in auxillary measurements. A straightforward approach is to go ahead and use the Garwood intervals at the point estimators for the nuisance parameters. Then the errors in the nuisance parameters can be regarded as errors in the model, or “systematic errors”, with the effect estimated by varying these parameters, i.e., by varying the model. But there are other options and the choice can depend on the particular circumstance, and such a discussion goes well beyond the scope of this paper. Simulations may also be required to justify a chosen approach.